# Diffusion Model as a Noise-Aware Latent Reward Model for Step-Level Preference Optimization

### v2

@@ Abstract
Preference optimization for diffusion models aims to align them with human preferences for images. Previous methods typically leverage Vision-Language Models (VLMs) as pixel-level reward models to approximate human preferences. However, when used for step-level preference optimization, these models face challenges in handling noisy images of different timesteps and require complex transformations into pixel space. In this work, we demonstrate that diffusion models are inherently well-suited for step-level reward modeling in the latent space, as they can naturally extract features from noisy latent images. Accordingly, we propose the Latent Reward Model (LRM), which repurposes components of diffusion models to predict preferences of latent images at various timesteps. Building on LRM, we introduce Latent Preference Optimization (LPO), a method designed for step-level preference optimization directly in the latent space. Experimental results indicate that LPO not only significantly enhances performance in aligning diffusion models with general, aesthetic, and text-image alignment preferences, but also achieves 2.5-28$\times$ training speedup compared to existing preference optimization methods. Our code and models are available at https://github.com/Kwai-Kolors/LPO.
Diffusion Model, Preference Optimization, Reward Model, Image Generation
[…]
@@ 1 Introduction
Diffusion models (Esser et al., 2024) have achieved significant success in the domain of text-to-image generation. Inspired by advancements in preference optimization (Touvron et al., 2023; Dubey et al., 2024) for Large Language Models (LLMs), several methods are proposed to align the diffusion model with human preferences. Diffusion-DPO (Wallace et al., 2024) extends Direct Preference Optimization (DPO) (Rafailov et al., 2023) to diffusion models, leveraging human-annotated data for training without requiring a reward model. However, the reliance on offline sampling introduces a distribution discrepancy between the preference data and the diffusion model, resulting in reduced optimization efficiency and effectiveness. DDPO (Black et al., 2024) and D3PO (Yang et al., 2024) employ online sampling and reward models for preference assessment. These methods typically assume that the preference order or reward value remains constant across all timesteps; however, this assumption may not hold in practice. To address this limitation, Step-by-step Preference Optimization (SPO) (Liang et al., 2024) introduces the Step-aware Preference Model (SPM) to predict preferences for different steps. Consequently, SPO can align models with aesthetic preference through step-level online sampling and preference assignment.
Figure 1: The comparison between the pixel-level reward model (a) and latent reward model (b) in step-level preference optimization. DMO denotes the diffusion model to be optimized.
Figure 2: Qualitative comparison among different preference optimization methods based on SDXL (Podell et al., 2023). LPO excels in both aesthetics and text-image alignment, resulting in improved overall quality. Larger versions are provided in Fig. 12 and Fig. 13.
@@
The reward models used in the aforementioned methods are mainly implemented by fine-tuning Vision-Language Models (VLMs) like CLIP (Radford et al., 2021) on preference datasets. We designate these models as Pixel-level Reward Models (PRMs) since they exclusively accept pixel-level image inputs. When applied to step-level preference optimization, PRMs encounter several common challenges. (1) Complex Transformation: At each timestep $t$, they necessitate additional processes of diffusion denoising and VAE (Kingma, 2013) decoding to transform noisy latent images $x_{t}$ into clean ones $\hat{x}_{0,t}$ and pixel-level images $I_{t}$. This results in an overly lengthy inference process, as illustrated in Fig. 1 (a) and Fig. 3 (a). (2) High-Noise Incompatibility: At large timesteps characterized by high noise, the predicted images $I_{t}$ are significantly blurred, leading to a severe distribution shift from the training data of VLMs, i.e. clear images. Consequently, the predictions of PRMs are unreliable at large timesteps. (3) Timestep Insensitivity: PRMs exhibit limited sensitivity to variations across different timesteps since they usually do not include timesteps as input, complicating the understanding of how different timesteps influence image assessment. These issues hinder the effectiveness of PRMs for step-level reward modeling.
@@
Is there a reward model that can naturally capture human preferences directly in the latent space while being aware of timesteps and compatible with high noise? In this paper, we find that the pre-trained diffusion model for text-to-image generation is an ideal choice because it exhibits several favorable characteristics, as listed in Fig. 1 (b). (1) It possesses inherent text-image alignment capabilities, owing to the pre-training on large-scale text-image pairs. (2) It can directly accept noisy latent images $x_{t}$ without requiring additional diffusion forwarding and VAE decoding. (3) It is high-noise-compatible since it can extract features from $x_{t}$ with various noise intensities, as is done during pre-training. (4) It naturally exhibits a strong sensitivity to timesteps, enabling it to effectively grasp the model’s attention at different timesteps. These pre-trained abilities make the diffusion model particularly suitable for step-level reward modeling.
@@
Based on the above insights, we propose using the diffusion model as a noise-aware Latent Reward Model (LRM). For noisy latent images $x_{t}$, LRM utilizes visual features from the UNet (Ronneberger et al., 2015) and textual features from the text encoder to predict step-level preference labels. We further introduce the Visual Feature Enhancement (VFE) module to enhance LRM’s focus on the text-image alignment. To address the inconsistent preference issue in LRM’s training data, we propose the Multi-Preference Consistent Filtering (MPCF) strategy to ensure that winning images consistently outperform losing ones across multiple preference dimensions, thus enabling LRM to effectively learn human preferences. Finally, we employ LRM for step-level preference optimization, leading to a simple yet effective method termed Latent Preference Optimization (LPO), where all steps are conducted within the latent space of diffusion models. Extensive experiments demonstrate that LPO significantly improves the image quality of various diffusion models and consistently outperforms existing DPO and SPO methods across the general, aesthetic, and alignment preferences, as indicated in Fig. 2. Moreover, LPO exhibits remarkable training efficiency, achieving a speedup of 10-28$\times$ over Diffuison-DPO and 2.5-3.5$\times$ over SPO.
@@ 2 Related Work
Reward Models for Human Preferences. Evaluating text-to-image generative models is a challenging problem. Several methods leverage Vision-Language Models (VLMs) to assess the alignment of generated images with human preferences. PickScore (Kirstain et al., 2023), HPSv2 (Wu et al., 2023), and ImageReward (Xu et al., 2023) aim to predict general preference by fine-tuning CLIP (Radford et al., 2021) or BLIP (Li et al., 2022) on preference datasets. Zhang et al. (2024) proposes CLIP-based MPS to capture multiple preference dimensions. Black et al. (2024) and Jiang et al. (2024) utilize LLaVA (Liu et al., 2023) and BLIP, respectively, to improve the text-image alignment of generative models. Liang et al. (2024) introduces SPM, based on PickScore, to predict step-level preference labels during the denoising process. These reward models, limited to accepting pixel-level images, often face problems like distribution shift and cumbersome inference. In contrast, our proposed LRM can effectively address these issues. To the best of our knowledge, we are the first to employ diffusion models themselves for reward modeling.
Preference Optimization for Diffusion Models. Motivated by improvements in Reinforcement Learning from Human Feedback (RLHF) in LLMs (Touvron et al., 2023; Dubey et al., 2024), several preference optimization approaches for diffusion models have been proposed. Differentiable reward fine-tuning methods (Clark et al., 2024; Wu et al., 2025) directly adjust diffusion models to maximize the reward of generated images. However, they are susceptible to reward hacking issues and require gradient backpropagation through multiple denoising steps, significantly increasing the computation burden. DPOK (Fan et al., 2023) and DDPO (Black et al., 2024) formulate the denoising process as a Markov decision process and employ Reinforcement Learning (RL) techniques to align models with specific preferences. Nonetheless, they exhibit inferior performance on open vocabulary sets. Diffusion-DPO (Wallace et al., 2024) and D3PO (Yang et al., 2024) apply DPO (Rafailov et al., 2023) in LLMs to diffusion models, yielding better performance than the aforementioned RL-based methods. These methods optimize the trajectory-level preference; however, the preference order in intermediate steps may not be the same as that of the final generated images. In view of this, SPO (Liang et al., 2024) proposes a step-level preference optimization method. In this paper, our proposed LPO optimizes models in the latent space by employing the diffusion model as a powerful and cost-effective reward model, demonstrating significant effectiveness and efficiency.
3 Background
Figure 3: The training pipelines of PRM (a) and LRM (b) for step-level reward modeling, and LPO (c). Compared to PRM, LRM can directly process noisy latent images $x_{t}$. Therefore, LPO is able to perform sampling and training within the latent space.
@@ 3.1 Latent Diffusion Models
The forward process of diffusion models gradually adds random Gaussian noise to clean latent images $x_{0}$ to obtain noisy latent images $x_{t}$, in the manner of a Markov chain:
$q(x_{t}|x_{t-1})=\mathcal{N}(\sqrt{\alpha_{t}}x_{t-1},(1-\alpha_{t})\mathbf{I}),$ (1)
where $t$ is the timestep and $\mathcal{N}(\mu,\Sigma)$ denotes the Gaussian distribution. $\alpha_{t}=1-\beta_{t}$, where $\beta_{t}$ is a pre-defined time-dependent variance schedule. $\mathbf{I}$ denotes the identity matrix. The backward process aims to denoise $x_{t}$, which can be formulated as follows according to DDIM (Song et al., 2021):
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{t},\sigma_{t}^{2}\epsilon_{t}),\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\$ (2)
$\mu_{t}=\sqrt{\bar{\alpha}_{t-1}}\hat{x}_{0,t}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta}^{(t)},\$ (3)
$\hat{x}_{0,t}=(\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta}^{(t)}}{\sqrt{\bar{\alpha}_{t}}}),\$ (4)
$\sigma_{t}=\eta\sqrt{(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})}\sqrt{1-\alpha_{t}},$ (5)
where $\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{t}$ and $\eta\in[0,1]$ is a hyperparameter to adjust the standard deviation $\sigma_{t}$. $\epsilon_{\theta}^{(t)}$ and $\hat{x}_{0,t}$ denote the noise and clean latent images predicted by diffusion models with parameters $\theta$ at timestep $t$.
@@ 3.2 Preference Optimization for Diffusion Models
Given winning latent images $x_{0}^{w}$, losing ones $x_{0}^{l}$, and condition $c$, Diffusion-DPO (Wallace et al., 2024) and D3PO (Yang et al., 2024) propagate the preference order of $(x_{0}^{w},x_{0}^{l})$ to all denoising steps, thereby freely generating intermediate preference pairs $(x_{t}^{w},x_{t}^{l})$. They encourage models $p_{\theta}$ to generate $x_{t}^{w}$ rather than $x_{t}^{l}$ by minimizing
$L_{DPO}=-\mathbb{E}_{\begin{subarray}{c}x^{w}_{t,t+1}\sim p_{\theta}(x^{w}_{t,t+1}|x^{w}_{0},c),\\ x^{l}_{t,t+1}\sim p_{\theta}(x^{l}_{t,t+1}|x^{l}_{0},c)\end{subarray}}\biggl{[}\log\sigma\biggl{(}\\ \beta\log\frac{p_{\theta}(x_{t}^{w}|x_{t+1}^{w},c)}{p_{ref}(x_{t}^{w}|x_{t+1}^{w},c)}-\beta\log\frac{p_{\theta}(x_{t}^{l}|x_{t+1}^{l},c)}{p_{ref}(x_{t}^{l}|x_{t+1}^{l},c)}\biggr{)}\biggr{]},$ (6)
where $\beta$ is a regularization hyperparameter and $p_{ref}$ denotes the reference model, fixed to the initial value of $p_{\theta}$.
However, the preference order along the denoising process is not always identical. Therefore, SPO (Liang et al., 2024) proposes SPM to predict preferences for intermediate steps, and sample the win-lose pairs $(x^{w}_{t},x^{l}_{t})$ from the same $x_{t+1}$ to make the two paths comparable. The optimization objective of SPO is to minimize
$L_{SPO}=-\mathbb{E}_{x^{w}_{t},x^{l}_{t}\sim p_{\theta}(x_{t}|x_{t+1},c)}\biggl{[}\log\sigma\biggl{(}\\ \beta\log\frac{p_{\theta}(x_{t}^{w}|x_{t+1},c)}{p_{ref}(x_{t}^{w}|x_{t+1},c)}-\beta\log\frac{p_{\theta}(x_{t}^{l}|x_{t+1},c)}{p_{ref}(x_{t}^{l}|x_{t+1},c)}\biggr{)}\biggr{]}.$ (7)
As a Pixel-level Reward Model (PRM), SPM faces some common issues discussed in Sec. 1. First, as illustrated in Fig. 3 (a), it requires the forward of the diffusion model and VAE decoder. Second, it struggles to process highly blurred $I_{t}$ when $t$ is large. Third, although the AdaLN block (Peebles & Xie, 2023) is introduced in SPM to improve its sensitivity to timesteps, insufficient pre-training makes it challenging to understand the focus of different timesteps on image generation. These issues jointly diminish the effectiveness of SPM in step-level preference optimization.
4 Methodology
To address the limitations of PRMs, we propose the Latent Reward Model (LRM), which leverages diffusion models for reward modeling. To facilitate its training, we analyze the preference dataset and propose a Multi-Preference Consistent Filtering (MPCF) strategy. Finally, we introduce Latent Preference Optimization (LPO), a method that optimizes diffusion models in the latent space using LRM.

### v3

@@ Abstract
Preference optimization for diffusion models aims to align them with human preferences for images. Previous methods typically use Vision-Language Models (VLMs) as pixel-level reward models to approximate human preferences. However, when used for step-level preference optimization, these models face challenges in handling noisy images of different timesteps and require complex transformations into pixel space. In this work, we show that pre-trained diffusion models are naturally suited for step-level reward modeling in the noisy latent space, as they are explicitly designed to process latent images at various noise levels. Accordingly, we propose the Latent Reward Model (LRM), which repurposes components of the diffusion model to predict preferences of latent images at arbitrary timesteps. Building on LRM, we introduce Latent Preference Optimization (LPO), a step-level preference optimization method conducted directly in the noisy latent space. Experimental results indicate that LPO significantly improves the model’s alignment with general, aesthetic, and text-image alignment preferences, while achieving a 2.5-28$\times$ training speedup over existing preference optimization methods. Our code and models are available at https://github.com/Kwai-Kolors/LPO.
@@ 1 Introduction
Diffusion models [1] have achieved significant success in the domain of text-to-image generation. Inspired by advancements in preference optimization [2, 3] for Large Language Models (LLMs), several methods are proposed to align the diffusion model with human preferences. Diffusion-DPO [4] extends Direct Preference Optimization (DPO) [5] to diffusion models, leveraging human-annotated data for training without requiring a reward model. However, the reliance on offline sampling introduces a distribution discrepancy between the preference data and the model, resulting in reduced optimization effectiveness. In contrast, DDPO [6] and D3PO [7] employ online sampling and reward models for preference assessment. Furthermore, to address the issue of preference order inconsistency at different timesteps, Step-by-step Preference Optimization (SPO) [8] introduces the Step-aware Preference Model (SPM) to predict step-wise preferences. Consequently, SPO can align models with aesthetic preference through step-level online sampling and preference assignment.
@@
Reward models are critical for the aforementioned methods and are mainly implemented by fine-tuning Vision-Language Models (VLMs) like CLIP [9] on preference datasets. We designate these models as Pixel-level Reward Models (PRMs) since they exclusively accept pixel-level image inputs. When applied to step-level preference optimization, PRMs encounter several common challenges. (1) Complex Transformation: At each timestep $t$, PRMs necessitate additional processes of diffusion denoising and VAE [10] decoding to transform noisy latent images $x_{t}$ into clean ones $\hat{x}_{0,t}$ and pixel-level images $I_{t}$. This results in an overly lengthy inference process, as illustrated in Fig. 2 (a) and Fig. 4 (a). (2) High-Noise Incompatibility: At large timesteps characterized by high-intensity noise, the predicted images $I_{t}$ are significantly blurred (as illustrated in Fig. 9), leading to a severe distribution shift from the training data of VLMs, i.e., clear images, thus making the predictions of PRMs at large timesteps unreliable. (3) Timestep Insensitivity: Since PRMs typically do not incorporate timesteps as input, it is challenging to understand the impact of different timesteps on image assessment. These issues hinder the effectiveness of PRMs for step-level reward modeling.
Figure 1: Qualitative comparison among different preference optimization methods based on SDXL [11]. LPO excels in both aesthetics and text-image alignment, resulting in improved overall quality. Larger versions with specific prompts are provided in Fig. 12 and Fig. 13.
Figure 2: The comparison between the pixel-level reward model (a) and the latent reward model (b). DMO denotes the diffusion model to be optimized.
@@
Is there a model that can naturally capture human preferences directly in the latent space while being aware of timesteps and compatible with high-intensity noise? We find that the pre-trained diffusion model for text-to-image generation is an ideal choice because it exhibits several favorable characteristics, as listed in Fig. 2 (b). (1) It possesses inherent text-image alignment capabilities, owing to the pre-training on large-scale text-image pairs. (2) It can directly process noisy latent images $x_{t}$ without requiring additional diffusion forwarding and VAE decoding. (3) It is high-noise-compatible since it can extract features from $x_{t}$ with various noise intensities, as is done during pre-training. (4) It naturally exhibits a strong sensitivity to the denoising timestep, enabling it to effectively grasp the model’s attention at different timesteps. These pre-trained abilities make the diffusion model particularly suitable for step-level reward modeling in the noisy latent space.
@@
Based on the above insights, we propose using the diffusion model as a noise-aware Latent Reward Model (LRM). Given noisy latent images $x_{t}$, LRM utilizes visual features from U-Net [12] or DiT [13] and textual features from the text encoder to predict step-level preference labels. A Visual Feature Enhancement (VFE) module is employed to enhance LRM’s focus on the text-image alignment. To address the inconsistent preference issue in LRM’s training data, we propose the Multi-Preference Consistent Filtering (MPCF) strategy to ensure that winning images consistently outperform losing ones across multiple dimensions. Finally, we employ LRM for step-level preference optimization, leading to a simple yet effective method termed Latent Preference Optimization (LPO), where all steps are conducted within the noisy latent space of diffusion models. Extensive experiments on SD1.5 [14] and SDXL [11] demonstrate that LPO substantially improves the quality of generated images and consistently outperforms existing DPO and SPO methods across the general, aesthetic, and alignment preferences, as indicated in Fig. 1. Meanwhile, LPO exhibits remarkable training efficiency, achieving a speedup of 10-28$\times$ over Diffuison-DPO [4] and 2.5-3.5$\times$ over SPO [8]. Furthermore, we explore a step-wise variant of GRPO [15] based on LRM and apply LPO to the DiT-based SD3 [1] model, demonstrating the generalization ability of LRM and LPO.
@@ 2 Related Work
Reward Models for Image Generation. Evaluating text-to-image generative models is a challenging problem. Several methods leverage VLMs to assess the alignment of images with human preferences. PickScore [16], HPSv2 [17], and ImageReward [18] aim to predict general preference by fine-tuning CLIP [9] or BLIP [19] on preference datasets. MPS [20] is proposed to capture multiple preference dimensions. Based on PickScore, SPM [8] is introduced to predict step-level preference labels during the denoising process. Recently, some works [21, 22, 23] employ more powerful VLMs to learn human preferences. However, these reward models, limited to accepting pixel-level images, often face problems like distribution shift and cumbersome inference when used in step-level preference optimization. In contrast, our proposed LRM can effectively mitigate these issues. To the best of our knowledge, we are the first to employ diffusion models themselves for reward modeling.
Preference Optimization for Diffusion Models. Motivated by improvements in Reinforcement Learning from Human Feedback (RLHF) in LLMs [2, 3], several optimization approaches for diffusion models have been proposed. Differentiable reward fine-tuning methods [24, 25, 26, 27] directly adjust diffusion models to maximize the reward of generated images. However, they are susceptible to reward hacking issues and require gradient backpropagation through multiple denoising steps. DPOK [28] and DDPO [6] formulate the denoising process as a Markov decision process and employ Reinforcement Learning (RL) techniques for preference alignment, but exhibit inferior performance on open vocabulary sets. Many works [4, 7, 29, 30, 31, 32] apply DPO [5] in LLMs to diffusion models, yielding better performance than the aforementioned RL-based methods. To mitigate the issue of inconsistent preference order across different timesteps, SPO [8] proposes a step-level preference optimization method. In addition, some works [33, 34] also explore training-free preference optimization methods. These methods typically perform optimization in pixel space, necessitating complex transformations and encountering distribution shift at large timesteps. On the contrary, LPO optimizes diffusion models directly in the noisy latent space using the LRM as a powerful and cost-effective reward model, demonstrating significant effectiveness and efficiency.
3 Preliminaries
@@ 3.1 Latent Diffusion Models
The forward process of diffusion models gradually adds random Gaussian noise to clean latent images $x_{0}$ to obtain noisy latent images $x_{t}$ at timestep $t$, in the manner of a Markov chain:
$q(x_{t}|x_{t-1})=\mathcal{N}(\sqrt{\alpha_{t}}x_{t-1},(1-\alpha_{t})\mathbf{I}),\quad\alpha_{t}=1-\beta_{t},$ (1)
where $\mathcal{N}(\mu,\Sigma)$ denotes the Gaussian distribution and $\beta_{t}$ is a pre-defined time-dependent variance schedule. $\mathbf{I}$ denotes the identity matrix. The backward process aims to denoise $x_{t}$, which can be formulated as follows according to DDIM [35]:
$p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(\mu_{t},\sigma_{t}^{2}\epsilon_{t}),\quad\epsilon_{t}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\$ (2)
$\mu_{t}=\sqrt{\bar{\alpha}_{t-1}}\hat{x}_{0,t}+\sqrt{1-\bar{\alpha}_{t-1}-\sigma_{t}^{2}}\cdot\epsilon_{\theta,t},\quad\hat{x}_{0,t}=(\frac{x_{t}-\sqrt{1-\bar{\alpha}_{t}}\epsilon_{\theta,t}}{\sqrt{\bar{\alpha}_{t}}}),\$ (3)
$\sigma_{t}=\eta\sqrt{(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})}\sqrt{1-\alpha_{t}},\quad\bar{\alpha}_{t}=\prod_{i=1}^{t}\alpha_{t},$ (4)
where $\eta\in[0,1]$ is a hyperparameter to adjust the standard deviation $\sigma_{t}$. $\epsilon_{\theta,t}$ and $\hat{x}_{0,t}$ denote the noise and clean latent images predicted by diffusion models with parameters $\theta$ at timestep $t$.
@@ 3.2 Preference Optimization for Diffusion Models
Given winning latent images $x_{0}^{w}$, losing ones $x_{0}^{l}$, and condition $c$, Diffusion-DPO [4] propagates the preference order of $(x_{0}^{w},x_{0}^{l})$ to all denoising steps, thereby freely generating intermediate preference pairs $(x_{t}^{w},x_{t}^{l})$. It encourage models $p_{\theta}$ to generate $x_{t}^{w}$ rather than $x_{t}^{l}$ by minimizing
$L_{DPO}=-\mathbb{E}_{\begin{subarray}{c}x^{w}_{t,t+1}\sim p_{\theta}(x^{w}_{t,t+1}|x^{w}_{0},c),\\ x^{l}_{t,t+1}\sim p_{\theta}(x^{l}_{t,t+1}|x^{l}_{0},c)\end{subarray}}\left[\log\sigma\left(\beta\log\frac{p_{\theta}(x_{t}^{w}|x_{t+1}^{w},c)}{p_{ref}(x_{t}^{w}|x_{t+1}^{w},c)}-\beta\log\frac{p_{\theta}(x_{t}^{l}|x_{t+1}^{l},c)}{p_{ref}(x_{t}^{l}|x_{t+1}^{l},c)}\right)\right],$ (5)
where $\beta$ is a regularization hyperparameter and $p_{ref}$ denotes the reference model, fixed to the initial value of $p_{\theta}$. However, the preference order along the denoising process is not always identical. Therefore, SPO [8] proposes SPM to predict preferences for intermediate steps, and sample the win-lose pairs $(x^{w}_{t},x^{l}_{t})$ from the same $x_{t+1}$ to make the two paths comparable. Accordingly, the optimization objective of SPO is reformulated as minimizing
$L_{SPO}=-\mathbb{E}_{x^{w}_{t},x^{l}_{t}\sim p_{\theta}(x_{t}|x_{t+1},c)}\biggl{[}\log\sigma\biggl{(}\beta\log\frac{p_{\theta}(x_{t}^{w}|x_{t+1},c)}{p_{ref}(x_{t}^{w}|x_{t+1},c)}-\beta\log\frac{p_{\theta}(x_{t}^{l}|x_{t+1},c)}{p_{ref}(x_{t}^{l}|x_{t+1},c)}\biggr{)}\biggr{]}.$ (6)
As a Pixel-level Reward Model (PRM), SPM faces some common issues discussed in Sec. 1. First, as illustrated in Fig. 4 (a) and (c), it requires additional procedures including $\hat{x}_{0,t}$ prediction and VAE decoding. Second, it struggles to process highly blurred $I_{t}$ when $t$ is large (Fig. 9). Third, although the AdaLN block [13] is introduced in SPM to improve its sensitivity to timesteps, insufficient pre-training makes it challenging to understand the focus of different timesteps on image generation. These issues jointly diminish the effectiveness of SPM in step-level preference optimization.
4 Method
To address the limitations of PRMs, we propose the Latent Reward Model (LRM) in Sec. 4.1, leveraging diffusion models for reward modeling. To facilitate its training, we propose a Multi-Preference Consistent Filtering (MPCF) strategy in Sec. 4.2. Finally, the Latent Preference Optimization (LPO) is introduced in Sec. 4.3 to optimize diffusion models directly in the noisy latent space.